OpenStudy (anonymous):

Solve by using tables. Give each answer to at most two decimal places. –2x^2 – 4 = –8x 0.59, 3.41 –3.41, –0.59 1.17, 6.83 –1.41, 1.41

OpenStudy (jiteshmeghwal9):

\[-2x^2-4+8x=0\]\[-2x^2+8x-4=0\]\[-(8) \pm \sqrt{(8)^2-4(-2)(-4)}\over2(-2)\]

OpenStudy (jiteshmeghwal9):

solve this to gt ur answer :)

OpenStudy (anonymous):

I have tried to solve that and its not working out :/

OpenStudy (anonymous):

Enter the problem into your graphing calculator's Table program. http://puu.sh/1hM6X

OpenStudy (anonymous):

i dont have a graphing calc :/

OpenStudy (anonymous):

I got 2 plus square root 2 or 2 minus square root of 2. Given square root of 2 is about 1.4 that gives about 0.6 and 3.4 as results

OpenStudy (anonymous):

Then you must solve by hand. \[ –2x^2 – 4 = –8x\\ 2x^2-8x+4=0\\ x^2-4x+2=0\\ x=\frac{4\pm\sqrt{8}}{2}\\ \ \ =2\pm\sqrt2 \]

OpenStudy (anonymous):

Solve by using tables. Give each answer to at most two decimal places. –2x2 – 4 = –8x (1 point) (1 pt) 0.59, 3.41 (0 pts) –3.41, –0.59 (0 pts) 1.17, 6.83 (0 pts) –1.41, 1.41 1 /1 point